5. If sec 4A = cosec (A – 20°), where 4A is an acute angle, find the value of A.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
sec 4A = cosec (A – 20°) We know that sec 4A = cosec (90° – 4A) To find the value of A, substitute the above equation in the given problem cosec (90° – 4A) = cosec (A – 20°) Now, equate the angles 90° – 4A= A- 20° 110° = 5A A = 110°/ 5 = 22° Therefore, the value of A = 22°